OpenStudy (yamyam70):

TIme rate of Change Differential Calculus A spherical balloon is being filled with a rate of 15feet^3 per sec. compute the time rate of change of the surface area of the balloon at the instant when its volume is 120 ft^3?

OpenStudy (xapproachesinfinity):

first when v=120 we have 4/3 pi r^3=120 solving for r r^3=90/pi r=(90/pi)^1/3 then we go to volume rate of change dv/dt=4pi r^2 dr/dt you need to solve for dr /dt then plug in that to dA/dt =8rdr/dt

OpenStudy (xapproachesinfinity):

that will give the rate of change in surface area

OpenStudy (xapproachesinfinity):

need further assistant?

OpenStudy (yamyam70):